# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v2

@@ 3.3 Disentangled Inversion Technique
Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal. Increasing the guidance scale enhances editability but sacrifices reconstruction accuracy due to latent code deviation during editing.
In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the target, and the harmonized branch. This bifurcation is designed to optimize the performance of each branch separately. For the source branch, we implement a targeted correction mechanism. By reintegrating the distance $z^{*}_{t}-z^{src}_{t}$ into $z^{src}_{t}$, we directly mitigate the deviation of the pathway. This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio. On the other hand, the target branch and harmonic branch is left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio. The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio. We will further discuss this in the Section 5.4.2. The algorithm of Disentangled Inversion Technique has been outlined in Algorithm 2.
@@
4 BenchMark Construction
Despite recent significant advancements in music generation utilizing the latent diffusion model, music editing has received comparably less attention, especially zero-shot music editing. Hence, to systematically validate our proposed method as a plug-and-play strategy for editing models and compare our method with existing zero-shot music editing methods, as well as compensate for the absence of standardized performance criteria for inversion and editing techniques, we construct a benchmark dataset, named ZoME-Bench (Zero-shot Music Editing Benchmark). ZoME-Bench comprises 1,100 audio samples which are selected from MusicCaps, spanning ten distinct editing types such as add instrument, delete instrument, change genre, and so on. Each sample in this collection is accompanied by its corresponding source prompt, target prompt, human instruction, and the source audio. Additionally, annotations pertinent to attention control methodologies, such as blended words, are meticulously included to facilitate detailed evaluations. We attach the details about our benchmark in Appendix B.
@@ 5 Experiments
5.1 Implementation Details
We perform the inference of different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023b) models with 200 inference steps. The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA 4090Ti GPU for a fair comparison. We evaluate all methods in ZoMo-Bench across all editing types. Details can be found in Appendix C.
5.2 Evaluation Metrics
In order to evaluate the performance of our novel Disentangled Inversion Control framework, we employ a diverse set of five metrics that encompass three critical dimensions of audio quality and fidelity: Struture Distance,CLAP Score,LPAPS,MSE (Mean Squared Error)and FAD (Fréchet Audio Distance). All the metrics are introduced in detail in Appendix C.1.
@@ 5.3 Zero-shot Music Editing Results
SDEdit 25.87 0.22 12.18 9.82 0.40
DDPM-Friendly 22.36 0.21 14.08 8.49 0.41
DDIM Inversion 17.11 0.18 6.24 6.12 0.55
DIC 11.97 0.15 2.49 4.54 0.60
Table 1: Comparison with baselines on ZoMo-Bench about struture and content preservation and CLAP similiarity. DIC: our proposed Disentangled Inversion Control.
We present a comparative study of our Disentangled Inversion Control(DIC) against several established music editing techniques, namely SDEdit (Liu et al., ), DDIM Inversion Ho et al. , and DDPM-Friendly (Manor & Michaeli, 2024). Utilizing the ZoMo-Bench test set we released, we assess the quality of the generated audio samples. The assessment primarily concentrates on two key aspects: Essential Content Preservation and Editing Fidelity. The results are summarized in Table 1, leading to two main conclusions:
(1) Our Disentangled Inversion Control demonstrates substantial advancements in maintaining essential content and enhancing editing fidelity, in comparison to the baseline methods. (2) Despite DDIM Inversion achieving high CLAP similarity scores, significantly underperforms in preserving essential content when compared to our approach.
(a) Spider chart for editing quality.
(b) Spider chart for editing consistency.
Figure 3: A comprehensive performance evaluation on the ZoME-bench. We present spider charts of editing quality (CLAP Scores) and consistency (LPAPS) across 9 editing tasks for DDPM-Friendly, DDIM Inversion, and Disentangled Inversion Control methods.
To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing. The LPAPS and CLAP scores are depicted in Figure 3. The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks. (2) While methods like DDIM Inversion exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing fine-grained manipulations, as reflected in their inferior performance, especially in the "Delete Instrument" and "Change Melody" non-rigid editing type.
@@ 5.4 Ablation Study
The ablation studies presented in this section aim to validate the contributions of the Unified Attention Control and Refining Dual-path Inversion to our framework’s overall performance. Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in the Appendix D.
@@ 5.4.1 Ablation on Attention Control Methods
HAC 11.97 0.15 2.49 4.54 0.60
w/ CAC 10.24 0.19 2.75 4.76 0.56
w/ MSAC 9.92 0.18 2.50 4.66 0.58
w/ HB 11.75 0.16 2.59 4.67 0.59
Table 2: Ablation results about different control methods. HAC: Harmonized Attention Control.
To ascertain the impact of our attention control mechanisms, we perform ablations on the following configurations: Cross-attention Control Only (w/ CAC), Mutual Self-attention Control only (w/ MSAC), without Harmonic Branch (w/o HB), and any attention control (w/o Control). The findings, detailed in Table 2, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance. (2) The combined use of CAC and MSAC further improves results, but the absence of the harmonic branch yields suboptimal outcomes.
@@ 5.4.2 Ablation on Disentangled Inversion Technique
To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from the Table 3: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score. 2) The incorporation of target distance into target latent and harmonic branch leads to an improvement in structure and content integrity but the CLAP score sharply decreases. 3) Add the harmonic distance into the harmonic latent witness a distant decrease.
@@
$[{\bm{d}}_{src},{\bm{0}},{\bm{d}}_{har}]$ 37.51 0.27 27.2 14.23 0.28
$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.60
Table 3: Ablation results about disentangled inversion technique. [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance to the corresponding branch in Algorithm 2.
@@ 5.5 Qualitative Results
To complement our quantitative findings, we present a qualitative comparison in Figure 4. Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with the preservation of melodic content and harmonic structure. In contrast, our Disentangled Inversion Control demonstrates superior performance in precise music editing while preserving structural integrity. We provide additional qualitative results in the Appendix G for further examination.
@@ 6 Conclusion
In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts. Our retrospective analysis revealed that while recent methods like DDIM inversion have propelled zero-shot editing forward, they are not without their shortcomings. We identified that DDIM inversion is plagued by an accumulation of errors throughout the diffusion steps, which detracts from its overall effectiveness. Moreover, the absence of a fine-tuned attention control mechanism was found to impede the nuanced editing of music. To address these issues, we introduced the Disentangled Inversion framework, a novel approach that meticulously separates the source and target branches to amplify their individual strengths in preservation and editing. In tandem, we presented the Harmonized Attention Control technique, which enhances the preservation of essential content by integrating the attention maps from the original audio into the diffusion process. Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,000 samples across 10 varied editing categories—attested to the superiority of our methods.
@@ Appendix A Preliminaries And Analyses
This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied in diffusion models for text-guided audio synthesis. It further explores the challenges associated with these methods.
@@ A.1 Diffusion Models
Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, which corresponds to the given conditioning prompt. We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
@@
Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$. At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using pre-trained network for $T$ steps. To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM samling:
@@ A.3 Classifier-free Guidance
where $\varnothing$ = ("") is the embedding of a null text. This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{"}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
@@ B.1 General information
Here are details about our ZoME-Bench dataset (Zero-shot Music Editing Benchmark).This dataset contains 1,000 audio samples , which are selected from MusicCaps , and each is 10 seconds long with a sample rate of 16k.
We refactored their captions to express an edit and divide them into 10 parts , each representing a type of editing. Sample and details are shown in the following table
@@ B.2 Annotation Process
We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023). With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type "change melody" as examples):
"There is a description of a Piece of music ,Please judge whether the description has information of melody.If not , just answer "Flase" ,else change its melody properly into the opposite one ,just change the adjective and don’t replace any instrument! "blended_word" is [origin melody, changed melody] "emphasize" is [changed melody] "blended_word" and "emphasize" are tuples.
Question:(A mellow, passionate melody from a noisy electric guitar) Answer:("original_prompt": "A mellow, [passionate] melody from a noisy electric guitar", "editing_prompt": "A mellow, [soft] melody from a noisy electric guitar", "blended_word": ["passionate melody","soft melody"], "emphasize" :["soft melody"]) Question:(A recording of solo harp music with a dreamy, relaxing melody.)
Answer:("original_prompt": "A recording of solo harp music with a dreamy, [relaxing] melody.", "editing_prompt": "A recording of solo harp music with a dreamy, [nervous] melody.", "blended_word": ["relaxing melody","nervous melody"], "emphasize" :["nervous melody"])
Question:("A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
Answer:("original_prompt": "A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", "editing_prompt": "A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", "blended_word": ["harmonized flute melody","passionate flute melody"]), "emphasize" :["passionate flute melody"])
now we have Question({origin caption})
Answer(?)"
In the same way , instructions were appended by prompt as follows(take type "change melody" as examples):
"There is two description of different pieces of music divided by &,Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
@@
Answer: change dissonant melody into harmonic melody  Question: [’original prompt’] & [’editing prompt’]
Answer(?)"
Through this method accmopnied with rounds of manual review ,we guarantee the quality of this Benchmark.

### v3

@@ 3.4 Disentangled Inversion Technique
Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal. Increasing the guidance scale enhances editability, but sacrifices reconstruction accuracy due to latent code deviation during editing.
In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the harmonic, and the target branch. This decoupling is designed to unleash the capabilities of each branch separately. For the source branch, we implement a targeted correction mechanism. By reintegrating the distance $z^{*}_{t}-z^{src}_{t}$ into $z^{src}_{t}$, we directly mitigate the deviation of the pathway. This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio. On the other hand, the target branch and harmonic branch are left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio. The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio. We will further discuss this in the Section 4.2.2. The algorithm of Disentangled Inversion Technique has been outlined in Algorithm 2.
@@ 4 Experiments
Implementation Details. We infer different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023c) models with 200 inference steps. The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA A800 GPU for a fair comparison. We evaluate all methods in ZoMo-Bench across all editing types with fixed length. And we use the MusicDelta subset of the MelodyDB (Bittner et al., 2014) dataset for variable length comparisons, comprised of 34 musical excerpts in varying styles and in lengths ranging from 20 seconds to 5 minutes. Details can be found in Appendix C.
Evaluation Metrics. Our models employ a comprehensive evaluation using both objective and subjective metrics to assess essential content preservation, text-audio alignment fidelity, and audio quality. Objective metrics include Structure Distance (SD) (Ju et al., 2023), CLAP (Contrastive Language-Audio Pretraining) Score (Elizalde et al., 2023), LPAPS (Learned Perceptual Audio Patch Similarity) (Iashin & Rahtu, 2021; Paissan et al., 2023), and FAD (Fréchet Audio Distance) (Kilgour et al., 2018). The CLAP Score evaluates how well the edited audio aligns with the target prompt, while Structure Distance, LPAPS, and FAD, which have been adapted from image domain metrics, measure the similarity between the edited and source audio. Specifically, LPAPS quantifies the consistency of the edited audio in relation to the source audio, and the FAD metric assesses the distance between two distributions of audio signals. The Structure Distance measures the structural similarity between the edited and source audio. For the subjective evaluation, we conduct crowd-sourced human assessments using the Mean Opinion Score (MOS) to evaluate both editing fidelity (MOS-Q) and content preservation (MOS-P). We attach the details of all metrics in Appendix C.1.
@@ 4.1 Zero-shot Music Editing Results
AudioLDM 2 23.86 0.21 10.36 0.58 73.48 70.12
MusicGen 23.39 0.21 6.63 0.59 75.46 71.28
SDEdit 25.87 0.22 12.18 0.40 69.38 66.23
DDIM Inversion 22.52 0.21 9.51 0.49 73.10 3.38
MusicMagus 16.23 0.19 5.15 0.55 75.12 74.34
DDPM-Friendly 18.30 0.19 5.16 0.54 75.27 73.86
MEDIC 11.97 0.15 2.49 0.61 79.81 77.29
Table 1: Comparison with baselines on ZoME-Bench with fixed length about structure, content preservation, and CLAP similarity.
We present a comparative study of our Disentangled Inversion Control (DIC) against several established music generation and editing baselines, including AudioLDM 2 (Liu et al., 2023c), MusicGen (Copet et al., 2023), SDEdit (Liu et al., 2023a), DDIM Inversion Ho et al. (2020), MusicMagus (Zhang et al., 2024), and DDPM-Friendly (Manor & Michaeli, 2024). Utilizing the ZoMo-Bench test set we developed, we assess the quality of the generated audio samples, focusing on two key aspects: Content Preservation and Editing Fidelity. The results, summarized in Table 1, lead to the following conclusions: (1) Our method, MEDIC, significantly outperforms both generation-based and inversion-based models in terms of editing fidelity and audio similarity, demonstrating its effectiveness in addressing complex editing tasks. (2) Although DDPM-Friendly and MusicMagus show improvements in essential content preservation, they struggle to maintain high text-audio alignment compared to generation models. MEDIC effectively addresses accumulated errors, achieving high audio similarity and an impressive CLAP score, reflecting its superior editing fidelity. (3) MEDIC also excels in subjective metrics, achieving the best performance in audio quality and content preservation.
Variable Length Comparisons We further evaluate MEDIC against baseline methods in a variable length setting, with the results presented in Table 2. Our analysis leads to the following conclusions: (1) MEDIC outperforms all baselines across all metrics, demonstrating the effectiveness of our approach in variable-length scenarios. (2) Inversion-based baselines experience a significant degradation in CLAP score, while MEDIC maintains the highest CLAP score of 0.51 on the MusicDelta dataset.
Method Objective Metrics Subjective Metrics
[…]
AudioLDM 2 24.40 0.22 7.07 0.44 66.37 64.28
MusicGen 27.71 0.23 7.70 0.46 67.41 63.76
SDEdit 28.12 0.24 13.21 0.24 62.80 62.18
DDIM Inversion 23.5 0.21 10.12 0.27 65.94 65.73
MusicMagus 25.6 0.22 7.13 0.43 67.45 67.12
DDPM-Friendly 21.53 0.23 6.68 0.30 66.34 67.28
MEDIC 19.5 0.20 6.58 0.51 71.62 70.18
Table 2: Comparison with baselines in variable length setting.
(a) Spider chart for CLAP score comparisons .
(b) Spider chart for FAD score comparisons.
Figure 3: A comprehensive performance evaluation on the ZoME-bench. We present spider charts of CLAP scores and FAD scores across 10 editing tasks for DDPM-Friendly, DDIM Inversion, MusicMagus, and Disentangled Inversion Control methods.
Fine-grained Comparisons on ZoME-Bench To substantiate the robustness of our method, we delve into fine-grained comparisons across different types of editing. The FAD and CLAP scores are depicted in Figure 3. The insights gleaned from this analysis are as follows: (1) Across all editing types, our Disentangled Inversion Control surpasses other methods, demonstrating its prowess in handling both rigid and non-rigid editing tasks. (2) While baselines exhibit capabilities in handling certain rigid editing scenarios, they fall short in executing non-rigid manipulations, as reflected in their inferior performance, especially in the “Change Genre" and “Change Melody" non-rigid editing types.
@@ 4.2 Ablation Study
The ablation studies presented in this section aim to validate the contributions of the Harmonized Attention Control and Disentangled Inversion to our framework’s overall performance. Additionally, we examine the impact of the classifier-free guidance scale on editing outcomes, with a comprehensive analysis included in Appendix D.
@@ 4.2.1 Ablation on Attention Control Methods
HAC 11.97 0.15 2.49 0.61
w/o MSA Control 14.13 0.19 2.75 0.56
w/o CA Control 13.78 0.18 2.50 0.58
w/o Harmonic Branch 12.75 0.16 2.59 0.59
Table 3: Ablation results about different control methods. MSA: Mutual Self-Attention, CA: Cross Attention.
To validate the impact of our attention control mechanisms, we perform ablations on the following configurations: Remove Mutual Self Attention Control (w/o MSA Control), Remove Cross Attention Control (w/o CA Control), and without Harmonic Branch (w/o Harmonic Branch). The findings, detailed in Table 3, reveal that: (1) Both cross-attention and mutual self-attention controls individually enhance editing performance. (2) Although the naive combination of mutual self-attention control and cross-attention control improves the preservation and fidelity, it still yields sub-optimal outcomes due to the lack of a progressive harmonic branch. This demonstrates the effectiveness of harmonized attention control.
@@ 4.2.2 Ablation on Disentangled Inversion Technique
To demonstrate the soundness of Algorithm 2 and the effective disentanglement of triple branches, we can draw conclusions from Table 4: 1) Incorporating source distance into the target latent and harmonic branch, this leads to a decline in both preservation metrics and CLAP similarity score. 2) Incorporating target distance into the target branch and harmonic branch improves structural integrity and content preservation. However, this modification leads to a significant decrease in the CLAP score, indicating that while structural quality may improve, the alignment with the target prompt suffers. 3) Adding harmonic distance to the harmonic branch leads to a noticeable decline in performance. This finding showcases that introducing excessive deviations in the harmonic branch may adversely affect the overall audio quality and coherence.
@@
$[{\bm{d}}_{src},{\bm{d}}_{har},{\bm{0}}]$ 37.51 0.27 27.2 14.23 0.28
$[{\bm{d}}_{src},{\bm{0}},{\bm{0}}]$ 11.97 0.15 2.49 4.54 0.61
Table 4: Ablation results from the disentangled inversion technique. [,,][$\cdot$,$\cdot$,$\cdot$][ ⋅ , ⋅ , ⋅ ] denotes adding the distance (line 6 in Algorithm 2). MSE means the mean square error loss between the edited audio features and source audio features.
@@ 4.3 Qualitative Results
To complement our quantitative findings, we present a qualitative comparison in Figure 4. Methods such as SDEdit and inversion-based techniques often struggle to balance high editability with preserving melodic content and harmonic structure. In contrast, our Disentangled Inversion Control performs better in precise music editing while preserving structural integrity. We provide additional qualitative results in the Appendix G for further examination.
@@ 5 Conclusion
In this paper, we explored the burgeoning realm of text-guided diffusion models for audio generation, recognizing their potential to reshape source audio in alignment with specific textual prompts. We proposed the Disentangled Inversion Control to support both rigid and non-rigid editing tasks. Instead of a two-branch setting, we add an intermediate branch named the harmonic branch to progressively integrate harmonic and melodic information in music by cross-attention control and mutual self-attention control. To counteract the accumulated errors caused by DDIM inversion and CFG, we introduced a simple but effective method named disentangled inversion to separate the diffusion process into triple branches and eliminate the latent discrepancy distance in the source branch. Our comprehensive evaluations, conducted on the ZoME-Bench—a robust benchmark for music editing comprising 1,100 samples across 10 varied editing categories—attested to the superiority of our methods. And the experiment results on variable length test sets and ablation studies further validated the effectiveness and robustness of our method. We envisage that our work could serve as a basis for future zero-shot music editing studies.
@@ Appendix A Preliminaries And Analyses
This section introduces the foundational concepts of DDIM sampling and classifier-free guidance as applied to diffusion models for text-guided audio synthesis. It further explores the challenges associated with these methods.
@@ A.1 Diffusion Models
Text-guided diffusion models aim to map a random noise vector ${\bm{z}}_{t}$ and textual condition ${\bm{c}}$ to an output audio ${\bm{z}}_{0}$, corresponding to the given conditioning prompt. We train a denoiser network $\epsilon_{\theta}({\bm{z}}_{t},t,{\bm{c}})$ to predict the Gaussian noise $\epsilon\in{\mathcal{N}}(0,{\mathbf{I}})$ following the objective:
@@
Where noise is added to the sampled data ${\bm{z}}_{0}$ according to timestamp $t$. At inference, given a noise vector ${\bm{z}}_{T}$, the noise $i$ is gradually removed by sequentially predicting it using a pre-trained network for $T$ steps. To generate audios from given ${\bm{z}}_{T}$, we employ the deterministic DDIM sampling:
@@ A.3 Classifier-free Guidance
where $\varnothing$ = (“”) is the embedding of a null text. This further leads to another perturbation from ${\bm{z}}^{{}^{\prime}}_{t}$ to ${\bm{z}}^{{}^{\prime\prime}}$ due to the destruction of the DDIM process and causes error augmentation as demonstrated in Figure 1.
@@ B.1 General information
Here are the details of our ZoME-Bench dataset (Zero-shot Music Editing Benchmark). This dataset contains 1,000 audio samples, selected from MusicCaps, with each sample being 10 seconds long and having a sample rate of 16k.
We refactor the original captions to express specific edits and divide them into 10 parts, each representing a different type of editing. A sample and details are shown in the following table 5.
@@ B.2 Annotation Process
We rebuild our caption from captions for Musiccaps offered by  (Agostinelli et al., 2023). With the help of ChatGPT-4 (OpenAI, 2023), we rebuild the caption with prompt as follows(take type “change melody" as examples):
Description: “There is a description of a Piece of music, Please judge whether the description has information of melody. If not, just answer “Flase", else change its melody properly into the opposite one, just change the adjective and don’t replace any instrument! ",“blended_word" is [origin melody, changed melody], “emphasize" is [changed melody], “blended_word" and “emphasize" are tuples.
Question: (A mellow, passionate melody from a noisy electric guitar)
Answer:(“source_prompt": “A mellow, [passionate] melody from a noisy electric guitar", “editing_prompt": “A mellow, [soft] melody from a noisy electric guitar", “blended_word": [“passionate melody", “soft melody"], “emphasize": [“soft melody"])
Question: (A recording of solo harp music with a dreamy, relaxing melody.)
Answer: (“source_prompt": “A recording of solo harp music with a dreamy, [relaxing] melody.", “editing_prompt": “A recording of solo harp music with a dreamy, [nervous] melody.", “blended_word": [“relaxing melody",“nervous melody"], “emphasize" :[“nervous melody"])
Question: (“A vintage, emotional song with mellow harmonized flute melody and soft wooden percussions")
Answer: (“source_prompt": “A vintage, emotional song with [passionate] flute melody and soft wooden percussions.", “editing_prompt": “A vintage, emotional song with [harmonized] flute melody and soft wooden percussions.", “blended_word": [“harmonized flute melody",“passionate flute melody"]), “emphasize" :[“passionate flute melody"])
Now we have Question:({origin caption}), Answer(?)"
In the same way, instructions are appended by prompt as follows (take type “change melody" as examples):
Description: “There are two descriptions of different pieces of music divided by &, Please describe the difference you need to give me the results in the following format: Question: this instrumental song features a [relaxing] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo & this instrumental song features a [cheerful] melody with a country feel accompanied by a guitar piano simple percussion and bass in a slow tempo
@@
Answer: change dissonant melody into harmonic melody
Now we have Question: [‘source prompt’] & [‘editing prompt’], Answer(?)"
Through this method, supplemented by rounds of manual review, we ensure the quality of this benchmark.